Determine h0 and e using the following given parameters: μ = 0.02756 Pa, N = 30 rev/s, W = 2210 N (bearing load), r = 19mm, c = 0.038 mm, and l = 38 mm. Also, determine the co-efficient of friction, the torque to overcome friction, power loss to friction, total volumetric flow rate Q, side flow rate Qs, maximum film pressure and the location of maximum and terminating pressures.


1
Expert's answer
2022-02-01T23:03:01-0500

μ = 0.02756 Pa

N = 30 rev/s

W = 2210 N (bearing load)

r = 19mm

c = 0.038 mm

l = 38 mm

"\delta=\frac{\mu N}{P}(\frac{D}{C})^2"

"\delta=\frac{0.02756*30}{\frac{2210}{38*38*10^{-6}}}(\frac{38}{0.076})^2=0.135"

"\frac{L}{D}=\frac{38}{38}=1"

From tables "\frac{h_0}{C}=0.42"

"\implies h_0=0.42*0.038=0.01596mm"

Eccentricity ratio, "\epsilon =0.58"

"e=\epsilon C=0.58*0.038=0.02204mm"